A father has a son and a daughter. The father is five times old as the daughter and the son is 2 years older than the daughter. In 10 years, the sum of their three ages will be 81. How old are they now?
Answer by ankor@dixie-net.com(22740)

A father has a son and a daughter.
The father is five times old as the daughter and the son is 2 years older than the daughter.
In 10 years, the sum of their three ages will be 81.
How old are they now?
:
let f = father's age
let s = son's age
let d = daughter's age
:
Write an equation for each statement:
" The father is five times old as the daughter"
f = 5d
" and the son is 2 years older than the daughter."
s = d + 2
" In 10 years, the sum of their three ages will be 81. "
(f+10) + s+10) + (d+10) = 81
f + s + d + 30 = 81
f + s + d = 81 - 30
f + s + d = 51
Replace f with 5d, replace s with (d+2)
5d + (d+2) + d = 51
7d = 51 - 2
d = 49/7
d = 7 yrs is the daughter's age
'
I'll let you find the ages of Dad and son, check by finding their sum in 10 yrs.
:
"